OpenStudy (anonymous):

If an investment of $2500 grows to $4406.43 in nine years with interest compounded annually, what is the interest rate? (Round your answer to one decimal place.)

OpenStudy (tkhunny):

One Deposit. Level Interest Rate.\[$2,500.00(1+i)^{9}\;=\;$4,406.43\]Solve for i.

OpenStudy (anonymous):

@tkhunny how would you work this problem out? I tried to solve it and it didn't work as I thought it would have.

OpenStudy (tkhunny):

What did you try? I would divide by $2,500.00, first. Then it looks like a logarithm would get us the rest of the way.

OpenStudy (anonymous):

@tkhunny Yep I did that then I got (1+i)^9=1.76 . Then how would you use log into this?

OpenStudy (tkhunny):

Let's please use more decimal places than that, 1.762572 You need this logathm rule: \[\log(a^{b}) = b\cdot\log(a)\]

OpenStudy (tkhunny):

* logarithm